OpenStudy (anonymous):

Hello Laura! I need help with a math problem. The volume of concrete needed is 12 cubic yds. I know the with is 12 feet and the height is 4 inches. How can I find the length of this driveway?

OpenStudy (wolf1728):

Width is 4 yards height is 4/36 yards = 1/9 yards = .1111111111 yards Volume needed is 12 cubic yards so length * width * height = 12 cubic yards length * 4 * .1111111 = 12 length = 12 / .444444444 length = 27 yards

OpenStudy (anonymous):

Thank you! I just wanted to make sure that I had the correct answer. You made my day! Thx again